# Sample size and power calculations for causal inference in observational studies

### v3

@@ B.3 Calculation of overlap coefficient $\phi$ with fitted propensity scores
In real applications, researchers do not have access to data and thus cannot calculate $\phi$ with fitted propensity scores. Nevertheless, we provide a computation method here for interested readers. This method is more robust than directly estimating the density and can be useful for simulation studies.
@@
$\phi=\int_{0}^{1}\sqrt{f_{0}(u)f_{1}(u)}\,\mathrm{d}u=\frac{\int_{0}^{1}\sqrt{u(1-u)}\Pr(e(X)=u)\,\mathrm{d}u}{\sqrt{r(1-r)}}.$ (B.2)
@@
$\phi=\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}F(u)}{\sqrt{r(1-r)}}\approx\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}\widehat{F}(u)}{\sqrt{r(1-r)}}=\frac{1}{n}\cdot\frac{\sum_{i=1}^{n}\sqrt{\widehat{e}_{i}(1-\widehat{e}_{i})}}{\sqrt{r(1-r)}}.$ (B.3)
C Extension to weighted average treatment effect estimands
We extend the above methodology to a general class of weighted average treatment effect (WATE), which includes the ATE and several other common estimands as special cases. Averaging the CATE $\tau(x)$ over a target population, we obtain the average treatment effect of that population. Specifically, assume the observed sample is drawn from a population with covariates distribution $f(x)$, and let $f(x)h(x)$ denote the covariate distribution of the target population, where $h(x)$ is called a tilting function. Then we can represent the average treatment effect on the target population by a WATE
$\tau^{h}=\frac{\operatorname{\mathbb{E}}[\tau(x)h(x)]}{\operatorname{\mathbb{E}}[h(x)]}.$ (C.1)
When $h(x)=1$, $\tau^{h}$ reduces to the ATE; when $h(x)=e(x)$, $\tau^{h}$ becomes the average treatment effect for the treated (ATT); when $h(x)=e(x)(1-e(x))$, $\tau^{h}$ becomes the average treatment effect for the overlap population (ATO). We can use a unified weighting strategy to identify the WATE because $\tau^{h}=\operatorname{\mathbb{E}}[w_{1}(X)ZY-w_{0}(X)(1-Z)Y]$, where $w_{1}(x)=h(x)/e(x),w_{0}(x)=h(x)/(1-e(x))$ (Li et al.,, 2018). This corresponds to (i) the inverse probability weight, $(w_{1}=1/e(x),w_{0}=1/(1-e(x)))$, for the ATE; (ii) the ATT weight, $(w_{1}=1,w_{0}=e(x)/(1-e(x)))$, for the ATT; and (iii) the overlap weight (OW), $(w_{1}=1-e(x),w_{0}=e(x))$, for the ATO.
The Hájek estimator for the WATE is:
$\widehat{\tau}_{w}=\frac{\sum_{i}w_{1}(X_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(X_{i})Z_{i}}-\frac{\sum_{i}w_{0}(X_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(X_{i})(1-Z_{i})}.$ (C.2)
We can adapt the procedures in Section 3 to derive sample size calculation formulas for the ATT and ATO estimands. Specifically, the procedures to determine the propensity score and the outcome distributions remain the same, but the variance formula changes. Unlike the ATE, there is no deterministic order between the variance of the Hájek estimator with the estimated versus true propensity score for other WATE estimands (Yang and Ding,, 2018). For the same reasons as with the ATE stated in Section 2, we choose to proceed with the variance with the true $e$.
Below we derive the variance of the Hájek estimator for the WATE with the tilting function $h$:
$V_{w}\equiv\operatorname{\mathbb{V}}(\widehat{\tau}_{w})=\frac{1}{\mathbb{E}[h(X_{i})^{2}]}\mathbb{E}\left[\left\{\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}+\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}\right\}h(X_{i})^{2}\right],$ (C.3)
where $\xi_{z}=\mathbb{E}_{h}[Y_{i}(z)]=\mathbb{E}[h(X_{i})Y_{i}(z)]/\mathbb{E}[h(X_{i})]$; the proof is relegated to Appendix A.8. For the ATE, $h(X_{i})=1$, and $V_{w}$ reduces to $V$ as in Equation 3. For the ATT and ATO, $h(X_{i})$ is a function of the propensity score $e(X_{i})$ and hence $V_{w}$ is also fully determined by the joint distribution of $e(X_{i})$ and the potential outcomes $Y_{i}(z)$, which can be computed via (16). However, either the ATT or the ATO has a closed-form expression of variance, but they can be computed numerically.
We outline how to compute $V_{w}$ with the tilting function $h(X)$ given the parameters $(\mu_{e},\sigma_{e})$ and $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$. For the ATT and ATO, $h(X)$ is a function of the propensity score $e(X)$, whose distribution can be approximated by $W_{e}$. Therefore, we can also write the tilting function as $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})$ for ATT and $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})[1-\operatorname{\mathrm{e}xpit}(W_{e})]$ for ATO, such that $h(W_{e})$ has approximately the same distribution as $h(X)$. Calculating the denominator $\mathbb{E}[h(X)^{2}]=\mathbb{E}[h(W_{e})^{2}]$ where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ involves numerically (e.g. Simpson’s method) evaluating the following integral
$\int_{-\infty}^{\infty}h(u)^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u,$ (C.4)
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.5)
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.6)
Similarly as the case for ATE, the result does not depend on $E_{1}$ or $E_{0}$, and only depends on $(r,\phi,\rho^{2})$ if one assumes $S_{1}^{2}=S_{0}^{2}$ and $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$.
D Additional Simulation and RHC Study Results
D.1 Density plot of the estimated propensity scores in the RHC study.
Figure 6: Density plot of the estimated propensity scores by treatment arm in RHC.
D.2 Summary quantities in Simulation 4.1
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00
$\phi$ 1.00 0.98 0.93 0.87 0.83 0.81
$r$ 0.50 0.50 0.50 0.50 0.50 0.50
$E_{1}$1.73 -1.90 -2.03 -2.13 -2.17 -2.19
$E_{0}$2.74 -2.56 -2.43 -2.34 -2.30 -2.27
$S_{1}^{2}$ 20.02 20.30 20.41 20.46 20.47 20.48
$S_{0}^{2}$ 20.03 19.70 19.46 19.28 19.20 19.14
$\rho_{1}^{2}$ 0.02 0.04 0.04 0.04 0.04 0.04
$\rho_{0}^{2}$ 0.03 0.03 0.02 0.02 0.02 0.02
$S^{2}$ 20.02 20.00 19.93 19.87 19.83 19.81
$\rho^{2}$ 0.02 0.03 0.03 0.02 0.02 0.02
$R^{2}$ 0.20 0.20 0.19 0.19 0.19 0.19
Table 2: Summary quantities $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, pooled summary quantities $(S^{2},\rho^{2})$ and the regression $R^{2}$ under different values of $\kappa$ in Simulation 4.1.
D.3 Simulations with binary outcomes
We repeat the simulation design as in Section 4.1, and dichotomize the potential outcome $Y_{i}(z)$ into a binary variable $Y^{b}_{i}(z)=\mathbb{I}(Y_{i}(z)>-2)$. The observed outcome $Y^{b}_{i}=Z_{i}Y^{b}_{i}(1)+(1-Z_{i})Y^{b}_{i}(0)$. The true ATE on $Y^{b}$ is $\tau^{b}=\mathbb{E}[Y^{b}_{i}(1)-Y^{b}_{i}(0)]=0.089$.
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00
$\phi$ 1.00 0.98 0.93 0.87 0.84 0.81
$r$ 0.50 0.50 0.50 0.50 0.50 0.50
$E_{1}$ 0.53 0.51 0.50 0.49 0.48 0.48
$E_{0}$ 0.43 0.45 0.46 0.47 0.47 0.48
$S_{1}^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$S_{0}^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$\rho_{1}^{2}$ 0.00 0.03 0.03 0.03 0.02 0.02
$\rho_{0}^{2}$ 0.00 0.02 0.01 0.01 0.01 0.01
$S^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$\rho^{2}$ 0.00 0.02 0.02 0.01 0.01 0.01
$R^{2}$ 0.13 0.13 0.12 0.12 0.12 0.12
Table 3: Summary quantities $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, pooled summary quantities $(S^{2},\rho^{2})$ and the regression $R^{2}$ under different values of $\kappa$ in Simulation D.3.
@@ D.3 Simulations with binary outcomes
Assume we estimate $\tau^{b}$ using the Hájek estimator with $Y^{b}_{i}$. Dichotomizing the outcome does not affect the propensity scores, and hence $(r,\phi)$ remain the same as in that in Simulation 4.1. The remaining quantities $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ and the pooled quantities $(S^{2},\rho^{2})$ are shown in Table 3 in Supplementary Material C.3. We inversely solve for the sample size to reach a nominal power of 0.80, and show the results on Table 4.
@@
$\phi$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 992 0.81 0.85 977 0.80 0.83 977 0.80 0.83
0.98 1019 0.80 0.85 977 0.80 0.82 1016 0.80 0.85
0.93 1166 0.80 0.85 975 0.74 0.78 1157 0.80 0.85
0.88 1528 0.83 0.85 973 0.62 0.68 1489 0.80 0.85
0.84 1972 0.83 0.87 972 0.56 0.59 1914 0.82 0.86
0.81 2460 0.86 0.89 972 0.49 0.51 2342 0.83 0.87
Table 4: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens,, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004.
The results show that the proposed method yields sample sizes that lead to powers close to 0.80, and the results are similar to Shook-Sa and Hudgens, (2022). In contrast, the two-sample $z$-test continues to lead to vastly underestimated sample size as the degree of overlap decreases.
D.4 Simulations with ATT and ATO estimands
With the same simulation design and data generating process as in Section 4.1, we focus on the ATT and ATO estimands, the true value of which are both 1. For $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$, corresponding to $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, we calculate the sample size of the Hájek estimator (C.2) for the ATT and ATO, with estimated and true propensity scores, respectively. Then we inverse the process under each scenario to calculate the power; the results are shown in Table 5. In all scenarios, the estimated sample size is close to the true value 1000. As predicted from the general theory on the WATE (Li et al.,, 2018), the ATT estimator has large variance than the ATO estimator. Consequently, given the same setting, the ATT requires larger sample size to achieve the same level of power the ATO, and the discrepancy increases as the degree of overlap decreases.
@@ D.4 Simulations with ATT and ATO estimands
Table 5: Sample size (size) calculated by the proposed method to achieve the nominal power 0.8 under $r=0.5$ and various degrees of overlap in simulations. The power of the Hájek estimators is obtained from bootstrapping 10,000 random samples of the calculated size.
As noted in Supplementary Material C, the sample size for the ATT and ATO calculated using the estimated propensity scores is not guaranteed to be conservative. However, this simulation shows that (i) the calculated sample sizes under the true and estimated propensity scores are similar, and (ii) the sample size calculated for the ATO appears to be conservative if one expects to use the estimated propensity score in the study.
D.5 RHC study with ATT and ATO estimands

### v4

@@ B.3 Calculation of overlap coefficient $\phi$ with fitted propensity scores
In practice, researchers may not have access to data and thus cannot calculate $\phi$ with fitted propensity scores. Nevertheless, we provide a computation method for interested readers. This method is more robust than directly estimating the density and can be useful for simulation studies.
@@
$\phi=\int_{0}^{1}\sqrt{f_{0}(u)f_{1}(u)}\,\mathrm{d}u=\frac{\int_{0}^{1}\sqrt{u(1-u)}\Pr(e(X)=u)\,\mathrm{d}u}{\sqrt{r(1-r)}}.$ (77)
@@
$\phi=\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}F(u)}{\sqrt{r(1-r)}}\approx\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}\widehat{F}(u)}{\sqrt{r(1-r)}}=\frac{1}{n}\cdot\frac{\sum_{i=1}^{n}\sqrt{\widehat{e}_{i}(1-\widehat{e}_{i})}}{\sqrt{r(1-r)}}.$ (78)
Appendix C Additional Simulation Results
C.1 Simulations with binary outcomes
@@
We repeat the simulation design in Section 5.1, and dichotomize the potential outcome $Y_{i}(z)$ into a binary variable $Y^{b}_{i}(z)=\mathbb{I}(Y_{i}(z)>-2)$. The observed outcome $Y^{b}_{i}=Z_{i}Y^{b}_{i}(1)+(1-Z_{i})Y^{b}_{i}(0)$. The true ATE on $Y^{b}$ is $\tau^{b}=\mathbb{E}[Y^{b}_{i}(1)-Y^{b}_{i}(0)]=0.089$. Assume that we estimate $\tau^{b}$ using the Hájek estimator with $Y^{b}_{i}$. Dichotomizing $Y_{i}(z)$ does not affect the propensity scores, and hence $\phi$ remain unchanged from Simulation 5.1, while $\rho^{2},R^{2}$ changed. The pooled $S^{2}=r(1-r)=0.25$. We inversely solve the sample size to reach a nominal power of 0.80, and show the results in Table 3.
Table 3: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004. The treatment proportion $r$ is fixed at 0.5.
@@
$\phi$ $\rho^{2}$ $R^{2}$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 0.00 0.13 992 0.81 0.85 977 0.80 0.83 977 0.80 0.83
0.98 0.02 0.13 1019 0.80 0.85 977 0.80 0.82 1016 0.80 0.85
0.93 0.02 0.12 1166 0.80 0.85 975 0.74 0.78 1157 0.80 0.85
0.88 0.01 0.12 1528 0.83 0.85 973 0.62 0.68 1489 0.80 0.85
0.84 0.01 0.12 1972 0.83 0.87 972 0.56 0.59 1914 0.82 0.86
0.81 0.01 0.12 2460 0.86 0.89 972 0.49 0.51 2342 0.83 0.87
From Table 3, the proposed method yields sample sizes achieving powers near 0.80, performing similarly to Shook-Sa and Hudgens (2022). In contrast, the two-sample $z$-test continuously leads to vastly underestimated sample size as the degree of overlap decreases.
C.2 Synthetic simulations with ATT and ATO estimands
With the same simulation design and data generating process as in Section 5.1, we focus on the ATT and ATO estimands, the true value of which are both 1. For $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$, corresponding to $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, we calculate the sample size of the Hájek estimator (16) for the ATT and ATO, with estimated and true propensity scores, respectively. Then we inverse the process under each scenario to calculate the power of the Hájek estimators; the results are shown in Table 4. In all scenarios, the estimated sample size is close to the true value 1000. As predicted from the general theory on the WATE (Li et al., 2018a), the ATT estimator has large variance than the ATO estimator. Consequently, given the same setting, the ATT requires larger sample size to achieve the same level of power the ATO, and the discrepancy increases as the degree of overlap decreases.
Table 4: Sample size (size) calculated by the proposed method to achieve the nominal power 0.8 under various degrees of overlap. The power is obtained from bootstrapping 10,000 random samples of the calculated size.
@@ C.2 Synthetic simulations with ATT and ATO estimands
The sample sizes for ATT and ATO calculated with estimated propensity scores are not guaranteed to be conservative. However, this simulation shows that (i) the sample sizes calculated with the true and estimated propensity scores are similar, and (ii) the sample size for the ATO appears to be conservative if one uses the estimated propensity score in the study.
C.3 Synthetic simulations with heterogeneous treatment effect and ATT and ATO estimand
Table 5 extends the HTE simulation to the ATT and ATO estimands, using the true estimand-specific effect size $\widetilde{\tau}=\tau^{h}/S$. The ATO remains well-calibrated, with empirical power near or above the nominal 0.80, consistent with its known efficiency and stability. The ATT, however, exhibits substantial power loss as overlap worsens – a sharp departure from the homogeneous case in Table 4.
Table 5: Sample size (size) calculated by the proposed method to achieve the nominal power 0.8 under various degrees of overlap and HTE. The power is obtained from bootstrapping 10,000 random samples of the calculated size.
ATT Power ATO Power
$\phi$ $\tau^{h}$ Size True PS Est PS $\tau^{h}$ Size True PS Est PS
covariates-driven HTE
1.00 1.00 667 0.81 0.89 1.00 665 0.80 0.89
0.98 1.03 652 0.78 0.85 1.00 689 0.80 0.89
0.93 1.06 729 0.72 0.77 0.99 742 0.81 0.87
0.87 1.08 986 0.64 0.68 0.99 798 0.78 0.87
0.83 1.09 1317 0.56 0.59 0.99 828 0.77 0.85
0.81 1.09 1689 0.50 0.51 0.99 846 0.76 0.84
propensity score-driven HTE
0.98 1.19 457 0.77 0.82 0.99 653 0.80 0.87
0.93 1.34 418 0.70 0.72 0.97 726 0.80 0.86
0.87 1.45 492 0.61 0.64 0.95 802 0.78 0.84
0.83 1.50 617 0.59 0.57 0.95 840 0.76 0.83
0.81 1.53 764 0.52 0.54 0.94 863 0.75 0.81
As noted in Appendix C.2, the sample size for the ATT is not guaranteed to be conservative; under HTE, this issue is amplified. Since HTE inflates $\operatorname{\mathbb{V}}\{Y(1)\}$ while leaving $\operatorname{\mathbb{V}}\{Y(0)\}$ unchanged, and all observed outcomes in the ATT’s target population are $Y_{i}(1)$, the variance underestimation from the omitted $\operatorname{\mathbb{V}}\{\tau(X)\}$ term (Remark 1) is particularly consequential for the ATT. Practitioners targeting the ATT should exercise greater caution when effect heterogeneity is suspected.
C.4 RHC simulation with ATT and ATO estimands